### v1

# PointNetPGAP-SLC: A 3D LiDAR-based Place Recognition Approach with Segment-level Consistency Training for Mobile Robots in Horticulture

Abstract
This paper addresses robotic place recognition in horticultural environments using 3D-LiDAR technology and deep learning. Three main contributions are proposed: (i) a novel model called PointNetPGAP, which combines a global average pooling aggregator and a pairwise feature interaction aggregator; (ii) a Segment-Level Consistency (SLC) model, used only during training, with the goal of augmenting the contrastive loss with a context-specific training signal to enhance descriptors; and (iii) a novel dataset named HORTO-3DLM featuring sequences from orchards and strawberry plantations. The experimental evaluation, conducted on the new HORTO-3DLM dataset, compares PointNetPGAP at the sequence- and segment-level with state-of-the-art (SOTA) models, including OverlapTransformer, PointNetVLAD, and LOGG3D. Additionally, all models were trained and evaluated using the SLC. Empirical results obtained through a cross-validation evaluation protocol demonstrate the superiority of PointNetPGAP compared to existing SOTA models. PointNetPGAP emerges as the best model in retrieving the top-1 candidate, outperforming PointNetVLAD (the second-best model). Moreover, when comparing the impact of training with the SLC model, performance increased on four out of the five evaluated models, indicating that adding a context-specific signal to the contrastive loss leads to improved descriptors. The dataset and the code will be made publicly available at https://github.com/Cybonic/PointNetPGAP-SLC.git
Figure 1: Geolocation, 3D maps and recording setup of the HORTO-3DLM dataset. Four data sequences were recorded over two years at different locations in the UK, using a Husky mobile platform equipped with a 32-beam Velodyne sensor and ZED-F9P RTK GPS/GNSS system.
00footnotetext: 1T.Barros, L.Garrote, P.Conde, C.Premebida and U.J.Nunes are with the University of Coimbra, Institute of Systems and Robotics, Department of Electrical and Computer Engineering, Portugal. E-mails: {tiagobarros, garrote, pedro.conde,  cpremebida, urbano}@isr.uc.pt00footnotetext: 2 M.J.Coombes and C.Liu are with the Dept. of Aeronautical and Automotive Engineering, LUCAS Lab, Loughborough University, UK. E-mails:{M.J.Coombes, C.Liu5}@lboro.ac.uk
I INTRODUCTION
Autonomous robots have become more common in modern agriculture, where they have been used for various tasks such as the application of pesticides and fertilizers, infrastructure inspection (e.g., fences and irrigation systems), and fruit and vegetable harvesting [1]. Agricultural robots require precise and accurate localization systems to operate reliably. In horticultural environments, plantations may be located within polytunnels, greenhouses or in places with lack of GNSS coverage. Therefore, in such environments, a GNSS-based localization systems may not be a reliable localization source. Alternatively, most localization systems rely on map-based approaches, where Simultaneous Localization and Mapping (SLAM) can be used to generate or update maps [2].
As part of SLAM systems, or as a standalone global localization approach, 3D LiDAR-based place recognition has seen significant advancements, driven primarily by the autonomous driving community, with the goal of achieving long-term localization [3, 4]. Despite recent efforts to adapt these methods to natural [5] and agricultural environments [6], major challenges still remain largely unresolved.
This work addresses the 3D LiDAR place recognition problem for robots operating in horticultural environments. Horticultural crops present unique challenges due to varying planting arrangements, both between different plant species and within the same species. Factors such as inter-row spacing, plant arrangement and density affect the rows’ permeability to LiDAR scans, allowing beams to penetrate multiple rows. This phenomenon generates sparse and highly overlapping scans with poor geometries, which leads to inter-row descriptor ambiguity.
To cope with these challenges, we propose three key contributions:
(i) PointNetPGAP: A lightweight 3D LiDAR-based place recognition model that utilizes PointNet [7] for local feature extraction and aggregates the local features by combining a global average pooling with another aggregator that explores pairwise feature interactions into a single descriptor.
(ii) Segment-Level Consistency (SLC) model: A model introduced during training to address the inter-row descriptor ambiguity problem. Specifically, we split the crops in various segments, and use the SLC model to provide a training signal, in addition to the contrastive loss, that differentiates scans from distinct rows/segments, even if they are highly overlapped. The SLC model consists of a Multi-layer Perceptron (MLP) which predicts the segment class for each descriptor and computes the loss based on the corresponding ground-truth. By relying on the descriptors as inputs, the SLC model is easily integrated with any place recognition model that outputs descriptors.
(iii) Horticulture 3D Localization and Mapping (HORTO-3DLM) Dataset: A new dataset from horticultural environments comprising data from three orchards and one strawberry plantation using a 3D LiDAR sensor and RTK-GPS mounted on a mobile platform, as shown in Fig. 1. Although, in this work, this dataset is primarily employed for place recognition, it can also be utilized for other localization and mapping tasks.
We benchmark PointNetPGAP against state-of-the-art (SOTA) models, namely OverlapTransformer [8], PointNetVLAD [9], and LOGG3D-Net [10] on the HORTO-3DLM dataset using a cross-validation evaluation protocol. The results demonstrate that PointNetPGAP offers a more effective and efficient solution compared to the aforementioned SOTA models. Furthermore, we train all models with the SLC approach and report the results both at the sequence- and segment-level. Notably, all models except OverlapTransformer exhibit performance improvements, which is primarily driven by an increase in performance in segments with higher ambiguity.
The remainder of this paper is organized as follows. In Section II, we discuss related work. Section III outlines the proposed method. The experimental evaluation and the results are presented and discussed in Section IV, while the concluding remarks are provided in Section V.
Figure 2: Top: Retrieval framework using PointNetPGAP to generate global descriptors. The framework comprises the following key stages: T1) An input scan $P$ is processed by PointNet to extract local features $F$; T2) These local features are aggregated into a global descriptor $D$ by concatenating the global average representation of local features with pairwise feature interactions into a single representation; T3) The global descriptor queries the database to retrieve the top-$k$ most similar places. Bottom: Training scheme with the following stages: B1) A training tuple consisting of scans, including an anchor-positive pair and $m$ negatives, is input into the model, which generates the corresponding descriptors; B2) In the SLC model, the descriptors are input into an MLP that predicts the corresponding segment classes. The loss is computed using the Negative Log-likelihood loss based on the predictions and ground-truth segment labels; B3) The LazyTriplet loss is calculated based on the distances between the anchor-positive pair and the anchor-negatives; B4) Finally, both losses are combined using a weight $\alpha$.
II RELATED WORK
Place recognition has been the subject of much research over the past decade, particularly in recent years, with a focus on 3D-LiDAR approaches [3]. LiDAR sensory data has been recognized as a relevant modality to achieve robustness in place recognition, primarily due to its relative invariance to challenging variations such as different seasonal or lighting conditions, which are important in horticultural robotics.
II-A 3D LiDAR-based Place Recognition
In the context of 3D LiDAR-based methods, although hand-crafted descriptors are still being used [11], Deep Learning (DL) techniques have emerged as the most prevalent choice for place modeling [3]. DL-based place recognition models can be broadly split into two main components: a local feature extraction module that maps the input scan to a local feature space, and a global feature aggregation module that aggregates the local features into a global descriptor.
As for the feature extraction module, approaches can be categorized based on how the point clouds are processed: those that extract features directly from point clouds, such as PointNetVLAD [9], and those that resort to proxy representations, such as voxels [12, 13], polar coordinates [14], or depth range images [15, 8]. Recent works such as MinkLoc3D [13], LCDNet [2], and LOGG3D-Net [10] tend to primarily use voxel-based approaches, employing sparse convolutions such as PV-RCNN [16] or 4D Spatio-Temporal ConvNets [17], which have shown higher feature extraction capabilities in urban-like environments. In this work, we resort to PointNet [18] due to its efficiency and lower computational demands (as demonstrated in Section IV-F).
Regarding aggregation strategies, most 3D LiDAR place recognition frameworks rely on first-order aggregation approaches such as NetVLAD [9, 19, 2], which computes the residuals w.r.t. to feature clusters, generalized-mean pooling (GeM) [13] or global max pooling (MAC) [20] which operate directly on the feature space. Inspired by these works, we propose an aggregation approach that combines, in a single descriptor, a first-order and a second-order aggregation approach. Specifically, we combine the global average pooling aggregator with a pairwise feature interaction aggregator into a single descriptor. We compare our hypothesis with the SOTA models LOGG3D-Net [10], PointNetVLAD [9], and OverlapTransformer [8], and show through results on real-world datasets that the combination of these two statistical-based aggregators yields higher performance than each aggregator individually and also the SOTA models.
II-B Datasets
3D LiDAR-based place recognition methods have gained traction as robust and scalable global localization approaches, primarily driven by the autonomous vehicle community, which works mainly on road/city environments [3]. Works such as MinkLoc3D [13], PointNetVLAD [9], LCDNet [2], or LOGG3D-Net [10] are evaluated on urban environments, using available 3D LiDAR urban datasets such as KITTI odometry [21], MulRan [22], Oxford RobotCar [23, 9], or KITTI-360 [24].
Less studied is 3D LiDAR place recognition in field or agricultural environments. One possible reason for this is the lack of available datasets which, recently, has been partially compensated by datasets such as Wild-places [25], RELLIS-3D [26], ORFD [27], TreeScope [6], or VineLiDAR [28], showing the increasing interest in such environments. Thus, given the lack of real-world data, especially in horticulture, we propose a novel dataset HORTO-3DLM that includes 3D LiDAR data of semi-structured horticulture environments, particularly from fruit trees namely apples and strawberries. The goal of HORTO-3DLM is to contribute with real-world quality data suitable to place recognition/localization in the domains related to field and agricultural robotics.
III Proposed Approach
This section introduces the PointNetPGAP and the proposed training regime, where a traditional contrastive learning approach is combined with the proposed SLC model. Figure 2 outlines both PointNetPGAP in a retrieval framework (Top) and PointNetPGAP in the proposed training scheme (Bottom), with the combination of both the contrastive and the Negative Log-Likelihood (NLL) losses.
III-A Problem Formulation
A retrieval-based place recognition framework comprises two stages (as illustrated in Fig. 2): firstly (denoted by T1 and T2), a LiDAR scan $P\in\mathbb{R}^{n\times 3}$ with $n$ points is mapped to a descriptor $D\in\mathbb{R}^{d}$, such that $\Theta:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{d}$; then (the 2nd stage: T3), the descriptor $D$ queries a database for the $k$ nearest neighbors (KNN) based on a similarity metric. The $k$ retrieved candidates represent the top-$k$ candidates for potential revisited places. In this work, we focus on the first part (i.e., T1 and T2) and follow a retrieval framework (i.e., T3) as described in [15].
III-B PointNetPGAP
PointNetPGAP, here defined as a mapping function $\Theta$ (outlined in Fig. 2 by T1, T2), can be decomposed into two functions: $\Theta\equiv\varphi\circ\phi$, where $\phi:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{n\times c}$ maps the input scan $P\in\mathbb{R}^{n\times 3}$ to hyper-dimensional features $F\in\mathbb{R}^{n\times c}$; and $\varphi:\mathbb{R}^{n\times c}\rightarrow\mathbb{R}^{d}$ aggregates the local features $F$ into a global descriptor $D\in\mathbb{R}^{d}$ with $d$ dimensions.
III-B1 Local Feature Extraction (T1)
In this work, we follow the same approach as in [9] and implement $\phi$ based on PointNet [7], a DL-based network that operates directly on LiDAR scans. Hence, it receives a LiDAR scan $P\in\mathbb{R}^{n\times 3}$ with $n$ points and returns a local feature representation $F\in\mathbb{R}^{n\times c}$ with $c$ dimensions.
III-B2 Feature Aggregation (T2)
As for the aggregation (i.e., $\varphi$), this work proposes the fusion of two statistically-inspired aggregators: a global average representation of the local features, also known as global average pooling, and an aggregator based on pairwise feature interactions.
The global average representation, hereafter referred to as GAP, is computed as follows:
$\bar{F}=\frac{1}{n}\sum_{f_{j}\in F}f_{j},$ (1)
where $f_{j}\in\mathbb{R}^{c}$ represents the $j$-th feature point, and $\bar{F}\in\mathbb{R}^{c}$ represents the global average representation. As for the aggregator based on the pairwise feature interactions, it is computed using the outer product of $F$ as follows:
$G=\frac{1}{n}F\,F^{T},$ (2)
where $G\in\mathbb{R}^{c\times c}$ represents pairwise feature interactions of the local features, which is flattened to form a vector representation $G^{\prime}\in\mathbb{R}^{c^{2}}$. Finally, to create the final descriptor $D$, both representations are concatenated into a single vector $D^{*}=[G^{\prime},\bar{F}]\in\mathbb{R}^{c+c^{2}}$, which is fed to a full-connected layer $\text{FC}:\mathbb{R}^{c+c^{2}}\rightarrow\mathbb{R}^{d}$, and normalized using the $L_{2}$-norm:
$D=L_{2}\text{-norm}(\text{FC}(D^{*})).$ (3)
III-C Network Training
Training PointNetPGAP involves a traditional contrastive approach using the LazyTriplet loss [9], along with the proposed SLC model. The training process is illustrated at the bottom of Fig. 2 and is divided into four main stages, labeled B1, B2, B3, and B4.
III-C1 Training Data (B1)
Place recognition methods traditionally resort to contrastive learning techniques, which learn to distinguish between similar (positive pairs) and dissimilar (negative pairs) data samples. In place recognition for urban environments, positive pairs are defined only by a fixed radius, which is a sufficient condition given the scale of the trajectories and revisits in these environments. However, to perform place recognition in crops with narrow rows, the radius is a necessary but not sufficient condition. The search conditions must guarantee that the search occurs along the row and does not transgress into neighboring rows. Splitting the crops into unique segments and confining the search within the segment ensures the necessary conditions.
Hence, in this work, a positive pair refers to a pair of scans that must meet the following constraints: (1) both scans must be within a $r_{th}$ range (in meters); (2) both scans must be from different revisits, meaning that the positive scan cannot be the immediate previous scan of the anchor; and (3) both scans must be from the same segment (see Fig. 3). Scan pairs that do not meet these three constraints belong to the set of negatives.
As such, the model is trained with tuples of scans (that satisfy the aforementioned conditions) $\tau^{P}=(P^{A},P^{P},\{P^{N}_{j}\})$ where $P^{A}\in\mathbb{R}^{n\times 3}$ is the anchor, $P^{P}\in\mathbb{R}^{n\times 3}$ is the closest positive scan w.r.t. $P^{A}$ in the Euclidean space, while $\{P^{N}_{j}|P^{N}_{j}\in\mathbb{R}^{n\times 3}\}_{j=1}^{m}$ is a set of $m$ randomly selected negatives. The scan tuple $\tau^{P}$ is mapped to the descriptor space (as described in Section III-B), generating the corresponding descriptor tuple $\tau^{D}=(D^{A},D^{P},\{D^{N}_{j}\})$.
III-C2 Segment-level consistency model (B2)
The SLC model introduces an additional signal during training, which is more context-related than the one generated by the contrastive loss. The goal is to resolve ambiguities created by scans that are highly overlapped but belong to neighboring segments. The SLC model is framed as a classification problem, aiming to predict the segment class using the descriptors as inputs. The loss between the predictions and the corresponding ground-truth segment class $s\in[1,2,...,L]$, where $L$ is the number of segments in a sequence (as shown in Fig. 3), is computed based on the NLL Loss. Hence, given a training tuple of descriptors $\tau^{D}$, and the corresponding tuple of ground-truth segment classes $\tau^{S}=(s^{A},s^{P},\{s^{N}_{j}\})$, all descriptors $D_{i}\in\tau^{D}$ are fed to an MLP to predict the corresponding segment class, such that:
$\mathbf{p}_{i}=\text{MLP}(D_{i}),$ (4)
where $\mathbf{p}_{i}=[p_{1},p_{2},...,p_{L}]$ with $p_{k}\in[0,1]$ representing the segment prediction scores which are used, together with the respective ground-truth segment label $s_{i}\in\tau^{S}$, to compute the loss as follows:
$\mathcal{L}_{S}=\sum_{i}NLLLoss(\mathbf{p}_{i},s_{i}).$ (5)
III-C3 Contrastive Loss (B3)
As for the contrastive cost function, the LazyTriplet loss [9] is used. Thus, given a descriptor tuple $\tau^{D}=(D^{A},D^{P},\{D^{N}_{j}\})$ the loss is computed as follows:
$\mathcal{L}_{T}=\text{max}(d_{AP}-d_{AN}+m,0),$ (6)
where $d_{AP}=\|D^{A}-D^{P}\|_{2}$ is the Euclidean distance between the anchor and the positive, $d_{AN}=\min\limits_{j}(\|D^{A}-\{D^{N}_{j}\}\|_{2})$ is the Euclidean distance between the anchor and the hardest negative i.e., the negative that is the closest w.r.t. $D^{A}$ in the descriptor space, while $m$ designates a margin value.
III-C4 Final Loss (B4)
Finally, both $\mathcal{L}_{T}$ and $\mathcal{L}_{S}$ are combined as follows:
$\mathcal{L}=\alpha\mathcal{L}_{T}+(1-\alpha)\mathcal{L}_{S},$ (7)
where $\alpha$ is adjustable weight.
IV EXPERIMENTAL EVALUATION
To support the experiments under realistic and challenging cross-horticultural domains, we introduce a new dataset comprising four sequences from distinct horticultural crops, including detailed annotations for ground truth evaluation. This section outlines the evaluation protocol, implementation details, and training procedures, including preprocessing steps and the architecture of the proposed approach. Additionally, quantitative metrics and qualitative retrieval results are reported, showcasing the effectiveness of the approach. Furthermore, we conduct a runtime analysis to assess the computational efficiency of the PointNetPGAP.
Figure 3: 2D and 3D visualizations of the sequences’ paths. The 2D representation shows the individual segments, while the 3D representation outlines the overlapping paths.
IV-A The HORTO-3DLM Dataset
For real-world deployment, it is important to capture specific conditions that may induce unique challenges in agricultural environments. For instance, complex conditions may arise due to varying field layouts (row length, row width, open field, within tunnels, etc.), which can differ from plant type to plant type and even within the same plant species. Another aspect that may present challenges is the changing appearance of the same plantation due to plant growth and flowering stages.
Hence, we created a dataset called Horticulture 3D Localization and Mapping (HORTO-3DLM)***https://github.com/Cybonic/HORTO-3DLM.git, which was recorded in the United Kingdom over two years at different locations. The dataset includes data collected from four crops: two orchards (apples and cherries), and one strawberry plantation within polytunnels with a table-top growing system. All sequences incorporate various revisits.
The data collection process involved teleoperating a Clearpath Husky mobile robot equipped with a Velodyne VLP32 3D LiDAR (10Hz) and a ZED-F9P RTK-GPS (5Hz). Figure 1 showcases the geolocation, the 3D maps of the four sequences, and the recording setup mounted on the mobile robot, while Table I provides more detailed information regarding each sequence. The three sequences from orchards are denominated OJ22, ON22, and OJ23, while the sequence from the strawberry plantation is designated SJ23.
TABLE I: Summary of the four sequences. The various sequence names represent the first letter of plantation type, month and year of the recording. For example, an orchard sequence recorded in November 2022 is named ON22. The columns have the following meaning: name of the sequence (Sequence), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
Sequence M Y Total Scans N.Rs N.Segs D [m] Plantation Type
Orchards
ON22 Nov. 2022 7974 4 6 514 Apple (open)
OJ22 July 2022 4361 3 5 206 Apple (open)
OJ23 June 2023 7229 3 5 459 Cherry (open)
Strawberries
SJ23 June 2023 6389 3 5 742 tunnels
TABLE II: Training and testing statistics used in the experiments.
Cross-validation
Sequence Test Anchors Training Anchors Seq. Training Anchors
OJ22 1797 1605 100
OJ23 4801 1193 512
ON22 4993 1210 495
SJ23 3509 1107 598
IV-B Methodology for Place Recognition
From the proposed HORTO-3DLM dataset described in Section IV-A, each sequence was split into various segments (S1, S2, …, S6) as illustrated in Fig. 3. Besides the rows, we also included the extremities of the fields (i.e., S3 and S4) as segments. Thus, for each data point (LiDAR scan), there exists a corresponding segment class.
Following the protocol defined in Section III-C1, $r_{th}$ was set to 2 m for the training data, which means that a positive pair exists whenever an anchor-positive pair is within a range of 2 m. Additionally, from all anchor-positive pairs, only the anchors that were at least 0.5 m apart were selected for training. For evaluation purposes, $r_{th}$ was set to 10 m, and all anchors were used.
IV-C Evaluation Protocol
The evaluation of the proposed approach is based on a k-fold cross-validation protocol, with $k=4$, which consists of training on three sequences and testing on the sequence that was left out. This is repeated four times to obtain results on all sequences.
The dataset split for the cross-validation protocol is outlined in Table II. The entries have the following meaning: the first column (Sequence) identifies the sequence on which a given model was tested, meaning the model was trained on the remaining sequences; the second column (Test Anchors) refers to the number of anchors/loops with which the model was tested; the third column (Training Anchors) refers to the number of anchors/loops the model was trained with, corresponding to the sum of the training anchors of the remaining sequences; the fourth column (Seq. Training Anchors) refers to the number of training anchors each sequence contains individually.
As for the performance metrics, the results obtained from the proposed experiments are reported using Recall for the top-k retrieved candidates (i.e., Recall@k).
IV-D Training and Implementation Details
All models were trained and evaluated under the same conditions. The input scans were downsampled to 10k points, and randomly rotated around the z-axis. The parameters of PointNetPGAP were the following: we used the original architecture of PoinNet, with T-Net disabled, descriptor size was set to 256 dimensions ($d=256$), the local features were set to 16 dimensions ($c=16$), the MLP was configured with three layers with 256, 64, and 6 neurons, respectively, while the weight of the final loss was set to 0.5 ($\alpha=0.5$). The SOTA models were implemented based on the original code.
All models were trained with 200 epochs on an NVIDIA GeForce RTX 3090 GPU, using the closest positive and 20 negatives ($m=20$) for each anchor. The margin value of the contrastive loss was set to 0.5 ($m=0.5$), and the model parameters were optimized using the AdamW optimizer with a learning rate ($L_{r}$) of 0.0001 and a weight decay ($W_{d}$) of 0.0005. Moreover, all proposed experiments were conducted on Python 3.8, and PyTorch with CUDA 11.6.
IV-E Retrieval Evaluation and Discussion
This section presents and discusses the empirical results. Firstly, we conduct an ablation study on the proposed PointNetPGAP model, delineating the contribution of each aggregator to the performance. Secondly, we benchmark PointNetPGAP, both trained with and without the SLC model, against SOTA models: PointNetVLAD [9], LOGG3D-Net [10], and OverlapTransformer [8]. Each of these models adopts distinct strategies for handling scans: PointNetVLAD processes scans directly, OverlapTransformer processes scans as range images, and LOGG3D-Net relies on a voxel-based representation and network. The main results are presented in Table III and Fig. 4, while results at the segment-level are presented in Fig. 3. Furthermore, a qualitative assessment is outlined in Fig. 6. Finally, the efficiency of the models is evaluated through runtime analysis. Throughout this section, PointNetPGAP will be abbreviated as PNPGAP, and PNPGAP-SLC when trained without and with the SLC model, respectively.
IV-E1 Ablation Study
The results of the ablation study are reported in Table IV. For this study, we report the Recall@1 and Recall@1% of each aggregation approach separately and combined. The results indicate that combining both aggregators yields higher performance compared to each aggregator individually, improving performance by 0.6 percentage points (pp) for retrieving the Top-1% candidates and 4.7 pp for retrieving the Top-1 candidate. These results suggest that the real impact of combining these aggregation approaches is on modeling the closest neighbor.
TABLE III: Recall@1 performance of the models trained: (without SLC)/(with SLC). The bolded scores represent the highest value between trained without and with SLC, while the underling indicates the highest score in each column.
  Model OJ22 OJ23 ON22 SJ23 MEAN OverlapTransformer [8] 0.084/0.075-0.009 0.153/0.164+0.011 0.221/0.194-0.027 0.084/0.073-0.011 0.136/0.127-0.009 LOGG3D [10] 0.357/0.394+0.037 0.309/0.329+0.020 0.412/0.470+0.058 0.254/0.285+0.031 0.333/0.370+0.037 PointNetVLAD [9] 0.560/0.659+0.099 0.485/0.502+0.017 0.611/0.625+0.014 0.676/0.717+0.041 0.583/0.630+0.043 PNPGAP(ours) 0.873/0.880+0.007 0.643/0.688+0.045 0.685/0.734+0.049 0.678/0.712+0.034 0.719/0.754+0.034
OJ22 Refer to caption
OJ23 Refer to caption
ON22 Refer to caption
SJ23 Refer to caption
Figure 4: Retrieval performance for the top-25 candidates on the four sequences.
Figure 5: Recall@1 performance at the segment-level for the four sequences. For more details on the segments, please see Fig. 3.
TABLE IV: Ablation study of PNPGAP, where GAP represents the global average pooling aggregator and PFI represents the aggregator based on pairwise feature interactions. The bolded scores represent the highest values in each column.
  GAP PFI Recall@1 Recall@1% OJ22 OJ23 ON22 SJ24 MEAN OJ22 OJ23 ON22 SJ24 MEAN 0.819 0.589 0.687 0.714 0.701 0.928 0.909 0.962 0.975 0.944 0.827 0.613 0.685 0.713 0.710 0.962 0.925 0.950 0.988 0.956 0.880 0.688 0.734 0.723 0.756 0.943 0.938 0.985 0.983 0.962
IV-E2 Comparison to SOTA
PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD (the second-best model) by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the top-1 candidate. This is also evident in Fig. 6, where a qualitative assessment is shown on the paths, displaying the true positive predictions of each model when retrieving the top-1 candidate.
When analyzing the performance of retrieving the top-25 candidates (Fig. 4), LOGG3D-Net, which relies on a voxel-based network for feature extraction, exhibits the highest increase in performance. Meanwhile, OverlapTransformer, which relies on range images and on a CNN-based network for feature extraction, shows poor performance across all sequences. The inadequate performance of OverlapTransformer is also apparent in Fig. 6, where in some segments the model fails to predict any loops correctly. This suggests the unsuitability of this model for agricultural environments.
At the segment-level, most models perform best in segments 3 and 4, representing the extremities of the crops. As illustrated in Fig. 1, these segments are distinctive, while segments representing the rows exhibit higher similarities and thus create ambiguities among neighboring rows, resulting in lower retrieval performance.
Figure 6: True positive predictions for retrieving the top-1 candidate.
IV-E3 Training with the SLC Model
The results when adding the SLC model at training time (presented in Table III) indicate that this additional training signal improves the performance throughout all sequences and models (except OverlapTransformer), with an average performance increase of 3.7 pp for LOGG3D, 4.3 pp for PointNetVLAD, and 3.34 pp for the proposed PNPGAP model.
These findings can be further analyzed in more detail in Fig. 3, where a comparison between PNPGAP and PNPGAP-SLC is made at the segment-level. The results show that, while the performance in segments with high performance remains mostly unchanged, segments such as those corresponding to rows (i.e., 1, 2, and 5) increase performance, indicating that the SLC model produces better descriptors that are less susceptible to ambiguities.
IV-F Runtime Analysis
TABLE V: The average runtime per batch, where the batch size is 20 scans.
Param. [M] OJ22 [ms] OJ23 [ms] ON22 [ms] SJ23 [ms] MEAN [ms]
OverlapTransformer 48.2 12.5 11.5 11.4 11.5 11.7
LOGG3D 8.8 108.3 115.9 108.0 82.0 103.5
PointNetVLAD 19.8 23.3 22.1 22.1 22.4 22.5
PNPGAP (ours) 0.4 3.3 3.5 3.3 3.4 3.38
The processing capacity and model size are crucial factors for real-world implementation, particularly in resource-constrained robotics applications. Throughout the runtime experiment, the results presented in Table V were obtained using identical hardware specifications as detailed in Section IV-D: an AMD Ryzen 9 5900X 12-Core CPU with 64 GB RAM and a GeForce RTX 3090 24GB GPU. The reported outcomes specifically denote the duration, expressed in milliseconds (ms), that each model required for processing a batch of 20 scans. The retrieval duration was omitted from this experiment because it remains uniform across all models. The size of the models is reported through the count of parameters, denominated in millions (M). PNPGAP emerges as one of the smaller models, comprising 0.4 million parameters and demanding on average 4.4 ms per batch.
A comparative analysis against other SOTA models reveals that PNPGAP has approximately 5% of the size of LOGG3D-Net and is 31 times faster. Furthermore, in comparison to PointNetVLAD, PNPGAP is only 2% of its size and is 6 times faster. These outcomes highlight the computational efficiency of the proposed PNPGAP model.
V CONCLUSIONS
In this letter, we addressed the 3D LiDAR place recognition problem in horticultural settings by introducing a novel aggregation approach that fuses two statistic-inspired aggregators into a model called PointNetPGAP; an SLC model to handle inter-row descriptor ambiguities; and a novel dataset, HORTO-3DLM, which includes four sequences: three from orchards and one from a strawberry plantation.
The experimental evaluation was conducted on the HORTO-3DLM dataset, and the proposed PointNetPGAP model was benchmarked against SOTA models. An ablation study demonstrated the superiority of the combined aggregators, while the results at both the sequence and segment levels show the proposed model outperforming SOTA models while being more efficient.
Regarding the proposed training approach, all models were trained with and without the SLC model. The results showed that, except for one model, incorporating the SLC model during training led to an increase in performance. This improvement was primarily observed in segments with high inter-row similarity. This suggests that providing a context-specific training signal, as offered by the SLC model, may be crucial for effectively training place recognition models in environments where LiDAR scans overlap significantly, even when originating from different rows.
ACKNOWLEDGMENTS
This work has been supported by the project GreenBotics (ref. PTDC/EEI-ROB/2459/2021), funded by Fundação para a Ciência e a Tecnologia (FCT), Portugal, and by Innovate UK under the Grant No. 10072930 and 10073332. It was also partially supported by FCT through grant UIDB/00048/2020 and under the PhD grant with reference 2021.06492.BD. The authors would also like to thank Dr Charles Whitfield at NIAB East Malling for facilitating orchard data collection campaigns.

### v2

@@
 Abstract
-This paper addresses robotic place recognition in horticultural environments using 3D-LiDAR technology and deep learning.
-Three main contributions are proposed: (i) a novel model called PointNetPGAP, which combines a global average pooling aggregator and a pairwise feature interaction aggregator; (ii) a Segment-Level Consistency (SLC) model, used only during training, with the goal of augmenting the contrastive loss with a context-specific training signal to enhance descriptors; and (iii) a novel dataset named HORTO-3DLM featuring sequences from orchards and strawberry plantations.
-The experimental evaluation, conducted on the new HORTO-3DLM dataset, compares PointNetPGAP at the sequence- and segment-level with state-of-the-art (SOTA) models, including OverlapTransformer, PointNetVLAD, and LOGG3D.
-Additionally, all models were trained and evaluated using the SLC.
-Empirical results obtained through a cross-validation evaluation protocol demonstrate the superiority of PointNetPGAP compared to existing SOTA models.
-PointNetPGAP emerges as the best model in retrieving the top-1 candidate, outperforming PointNetVLAD (the second-best model).
-Moreover, when comparing the impact of training with the SLC model, performance increased on four out of the five evaluated models, indicating that adding a context-specific signal to the contrastive loss leads to improved descriptors.
+3D LiDAR-based place recognition remains largely underexplored in horticultural environments, which present unique challenges due to their semi-permeable nature to laser beams.
+This characteristic often results in highly similar LiDAR scans from adjacent rows, leading to descriptor ambiguity and, consequently, compromised retrieval performance.
+In this work, we address the challenges of 3D LiDAR place recognition in horticultural environments, particularly focusing on inter-row ambiguity by introducing three key contributions: (i) a novel model, PointNetPGAP, which combines the outputs of two statistically-inspired aggregators into a single descriptor; (ii) a Segment-Level Consistency (SLC) model, used exclusively during training to enhance descriptor robustness; and (iii) the HORTO-3DLM dataset, comprising LiDAR sequences from orchards and strawberry fields.
+Experimental evaluations conducted on the HORTO-3DLM and KITTI Odometry datasets demonstrate that PointNetPGAP outperforms state-of-the-art models, including OverlapTransformer and PointNetVLAD, particularly when the SLC model is applied.
+These results underscore the model’s superiority, especially in horticultural environments, by significantly improving retrieval performance in segments with higher ambiguity.
 The dataset and the code will be made publicly available at https://github.com/Cybonic/PointNetPGAP-SLC.git
@@
 I INTRODUCTION
-Autonomous robots have become more common in modern agriculture, where they have been used for various tasks such as the application of pesticides and fertilizers, infrastructure inspection (e.g., fences and irrigation systems), and fruit and vegetable harvesting [1].
-Agricultural robots require precise and accurate localization systems to operate reliably.
-In horticultural environments, plantations may be located within polytunnels, greenhouses or in places with lack of GNSS coverage.
-Therefore, in such environments, a GNSS-based localization systems may not be a reliable localization source.
-Alternatively, most localization systems rely on map-based approaches, where Simultaneous Localization and Mapping (SLAM) can be used to generate or update maps [2].
-As part of SLAM systems, or as a standalone global localization approach, 3D LiDAR-based place recognition has seen significant advancements, driven primarily by the autonomous driving community, with the goal of achieving long-term localization [3, 4].
-Despite recent efforts to adapt these methods to natural [5] and agricultural environments [6], major challenges still remain largely unresolved.
-This work addresses the 3D LiDAR place recognition problem for robots operating in horticultural environments.
-Horticultural crops present unique challenges due to varying planting arrangements, both between different plant species and within the same species.
-Factors such as inter-row spacing, plant arrangement and density affect the rows’ permeability to LiDAR scans, allowing beams to penetrate multiple rows.
-This phenomenon generates sparse and highly overlapping scans with poor geometries, which leads to inter-row descriptor ambiguity.
-To cope with these challenges, we propose three key contributions:
-(i) PointNetPGAP: A lightweight 3D LiDAR-based place recognition model that utilizes PointNet [7] for local feature extraction and aggregates the local features by combining a global average pooling with another aggregator that explores pairwise feature interactions into a single descriptor.
-(ii) Segment-Level Consistency (SLC) model: A model introduced during training to address the inter-row descriptor ambiguity problem.
-Specifically, we split the crops in various segments, and use the SLC model to provide a training signal, in addition to the contrastive loss, that differentiates scans from distinct rows/segments, even if they are highly overlapped.
-The SLC model consists of a Multi-layer Perceptron (MLP) which predicts the segment class for each descriptor and computes the loss based on the corresponding ground-truth.
-By relying on the descriptors as inputs, the SLC model is easily integrated with any place recognition model that outputs descriptors.
-(iii) Horticulture 3D Localization and Mapping (HORTO-3DLM) Dataset: A new dataset from horticultural environments comprising data from three orchards and one strawberry plantation using a 3D LiDAR sensor and RTK-GPS mounted on a mobile platform, as shown in Fig.
-1.
-Although, in this work, this dataset is primarily employed for place recognition, it can also be utilized for other localization and mapping tasks.
-We benchmark PointNetPGAP against state-of-the-art (SOTA) models, namely OverlapTransformer [8], PointNetVLAD [9], and LOGG3D-Net [10] on the HORTO-3DLM dataset using a cross-validation evaluation protocol.
-The results demonstrate that PointNetPGAP offers a more effective and efficient solution compared to the aforementioned SOTA models.
-Furthermore, we train all models with the SLC approach and report the results both at the sequence- and segment-level.
-Notably, all models except OverlapTransformer exhibit performance improvements, which is primarily driven by an increase in performance in segments with higher ambiguity.
-The remainder of this paper is organized as follows.
-In Section II, we discuss related work.
-Section III outlines the proposed method.
-The experimental evaluation and the results are presented and discussed in Section IV, while the concluding remarks are provided in Section V.
-Figure 2: Top: Retrieval framework using PointNetPGAP to generate global descriptors.
-The framework comprises the following key stages: T1) An input scan $P$ is processed by PointNet to extract local features $F$; T2) These local features are aggregated into a global descriptor $D$ by concatenating the global average representation of local features with pairwise feature interactions into a single representation; T3) The global descriptor queries the database to retrieve the top-$k$ most similar places.
-Bottom: Training scheme with the following stages: B1) A training tuple consisting of scans, including an anchor-positive pair and $m$ negatives, is input into the model, which generates the corresponding descriptors; B2) In the SLC model, the descriptors are input into an MLP that predicts the corresponding segment classes.
-The loss is computed using the Negative Log-likelihood loss based on the predictions and ground-truth segment labels; B3) The LazyTriplet loss is calculated based on the distances between the anchor-positive pair and the anchor-negatives; B4) Finally, both losses are combined using a weight $\alpha$.
+In recent years, significant advancements have been made in 3D LiDAR-based place recognition, driven primarily by the autonomous driving community with the aim of achieving long-term localization [1, 2].
+While initial efforts have been made to adapt these methods for agricultural environments [3, 4], several challenges unique to these settings remain largely unresolved.
+Unlike urban environments, where LiDAR scans capture well-defined geometries (e.g., corners, planes), horticultural environments produce scans with sparse and poorly defined features.
+The semi-permeable nature of these environments allows laser beams to traverse multiple row layers, resulting in scans that cover several rows.
+This phenomenon leads to sparse scans that overlap with scans from adjacent rows, thereby causing both intra- and inter-row descriptor ambiguity.
+This work addresses the challenges of 3D LiDAR place recognition in horticultural environments, particularly focusing on inter-row ambiguity, by introducing three key contributions: (i) PointNetPGAP: A lightweight model that combines PointNet for local feature extraction and integrates the outputs of two statistically-inspired aggregation approaches into a single descriptor.
+(ii) Segment-Level Consistency (SLC) Model: A training model designed to reduce inter-row ambiguity by differentiating scans from different segments.
+(iii) HORTO-3DLM Dataset: A novel dataset collected using a 3D LiDAR sensor, comprising data from three orchards and a strawberry plantation.
+We benchmark PointNetPGAP at both the sequence and segment levels against state-of-the-art (SOTA) models, specifically OverlapTransformer [5], PointNetVLAD [6], and LOGG3D-Net [7], using the HORTO-3DLM dataset within a cross-validation protocol.
+Additionally, we evaluate the cross-domain generalization capabilities using the KITTI Odometry dataset [8].
+The results highlight that the aggregation approach of PointNetPGAP is better suited to horticultural environments compared to existing SOTA models.
+Although its impact is reduced in urban settings, it nonetheless exhibits superior rotation invariance in urban environments compared to SOTA models.
+Moreover, when models are trained with the SLC approach, results at both the sequence and segment levels indicate improved , particularly by improving retrieval accuracy in segments with higher ambiguity.
+Figure 2: Top: The retrieval framework uses PointNetPGAP to generate global descriptors through the following stages: T1) PointNet processes an input scan $P$ to extract local features $F$; T2) These features are aggregated into a global descriptor $D$ by combining global average representation with pairwise feature interactions; T3) The global descriptor is then used to query the database and retrieve the top-$k$ most similar places.
+Bottom: The training scheme involves: B1) Inputting a training tuple with an anchor-positive pair and $m$ negatives to generate descriptors; B2) Feeding these descriptors into an MLP in the SLC model to predict segment classes, the SLC loss computed via Negative Log-likelihood using the predictions and the segment labels; B3) Calculating the LazyTriplet loss based on anchor-positive and anchor-negative distances; B4) Combining both losses with a weight $\alpha$.
 II RELATED WORK
-Place recognition has been the subject of much research over the past decade, particularly in recent years, with a focus on 3D-LiDAR approaches [3].
-LiDAR sensory data has been recognized as a relevant modality to achieve robustness in place recognition, primarily due to its relative invariance to challenging variations such as different seasonal or lighting conditions, which are important in horticultural robotics.
+Place recognition has been the focus of extensive research in recent years, particularly within the autonomous driving community [1], where Deep Learning-based approaches have garnered the most attention.
+One potential reason for the limited research in other domains may be the scarcity of publicly available datasets.
 II-A 3D LiDAR-based Place Recognition
@@
 Regarding aggregation strategies, most 3D LiDAR place recognition frameworks rely on first-order aggregation approaches such as NetVLAD [6, 18, 14], which computes the residuals w.r.t.
-to feature clusters, generalized-mean pooling (GeM) [13] or global max pooling (MAC) [20] which operate directly on the feature space.
+feature clusters, generalized-mean pooling (GeM) [11] or global max pooling (MAC) [19] which operate directly on the feature space.
 Inspired by these works, we propose an aggregation approach that combines, in a single descriptor, a first-order and a second-order aggregation approach.
@@
 One possible reason for this is the lack of available datasets which, recently, has been partially compensated by datasets such as Wild-places [23], RELLIS-3D [24], ORFD [25], TreeScope [4], or VineLiDAR [26], showing the increasing interest in such environments.
-Thus, given the lack of real-world data, especially in horticulture, we propose a novel dataset HORTO-3DLM that includes 3D LiDAR data of semi-structured horticulture environments, particularly from fruit trees namely apples and strawberries.
-The goal of HORTO-3DLM is to contribute with real-world quality data suitable to place recognition/localization in the domains related to field and agricultural robotics.
 III Proposed Approach
@@
 $\bar{F}=\frac{1}{n}\sum_{f_{j}\in F}f_{j},$ (1)
-where $f_{j}\in\mathbb{R}^{c}$ represents the $j$-th feature point, and $\bar{F}\in\mathbb{R}^{c}$ represents the global average representation.
-As for the aggregator based on the pairwise feature interactions, it is computed using the outer product of $F$ as follows:
-$G=\frac{1}{n}F\,F^{T},$ (2)
-where $G\in\mathbb{R}^{c\times c}$ represents pairwise feature interactions of the local features, which is flattened to form a vector representation $G^{\prime}\in\mathbb{R}^{c^{2}}$.
-Finally, to create the final descriptor $D$, both representations are concatenated into a single vector $D^{*}=[G^{\prime},\bar{F}]\in\mathbb{R}^{c+c^{2}}$, which is fed to a full-connected layer $\text{FC}:\mathbb{R}^{c+c^{2}}\rightarrow\mathbb{R}^{d}$, and normalized using the $L_{2}$-norm:
-$D=L_{2}\text{-norm}(\text{FC}(D^{*})).$ (3)
+where $f_{j}\in\mathbb{R}^{c}$ represents the $j$-th feature vector, and $\bar{F}\in\mathbb{R}^{c}$ represents the global average representation.
+As for the aggregator based on the pairwise feature interactions, it is computed as follows:
+$G=\frac{1}{n}F^{T}\,F,$ (2)
+where $G\in\mathbb{R}^{c\times c}$ represents pairwise feature interactions of the local features, which is flattened into a vector representation $G^{\prime}\in\mathbb{R}^{c^{2}}$.
+Finally, to create the final descriptor $D$, both representations are concatenated into a single vector $D^{*}=[G^{\prime},\bar{F}]\in\mathbb{R}^{c+c^{2}}$, which is fed to a full-connected layer $\text{FC}:\mathbb{R}^{c+c^{2}}\rightarrow\mathbb{R}^{d}$, and normalized using the $L_{2}$-norm: $D=L_{2}\text{-norm}(\text{FC}(D^{*}))$.
 III-C Network Training
@@
 The search conditions must guarantee that the search occurs along the row and does not transgress into neighboring rows.
-Splitting the crops into unique segments and confining the search within the segment ensures the necessary conditions.
+Dividing the crops into distinct segments and confining the search within each segment ensures the necessary conditions.
 Hence, in this work, a positive pair refers to a pair of scans that must meet the following constraints: (1) both scans must be within a $r_{th}$ range (in meters); (2) both scans must be from different revisits, meaning that the positive scan cannot be the immediate previous scan of the anchor; and (3) both scans must be from the same segment (see Fig.
@@
 The scan tuple $\tau^{P}$ is mapped to the descriptor space (as described in Section III-B), generating the corresponding descriptor tuple $\tau^{D}=(D^{A},D^{P},\{D^{N}_{j}\})$.
+Figure 3: 2D and 3D visualizations of the sequences’ paths. The 2D representation shows the individual segments, while the 3D representation outlines the overlapping paths.
 III-C2 Segment-level consistency model (B2)
@@
 The SLC model is framed as a classification problem, aiming to predict the segment class using the descriptors as inputs.
-The loss between the predictions and the corresponding ground-truth segment class $s\in[1,2,...,L]$, where $L$ is the number of segments in a sequence (as shown in Fig.
+The loss between the predictions and the corresponding ground-truth segment class $s\in\{1,2,...,L\}$, where $L$ is the number of segments in a sequence (as shown in Fig.
 3), is computed based on the NLL Loss.
@@
 $\mathbf{p}_{i}=\text{MLP}(D_{i}),$ (3)
-where $\mathbf{p}_{i}=[p_{1},p_{2},...,p_{L}]$ with $p_{k}\in[0,1]$ representing the segment prediction scores which are used, together with the respective ground-truth segment label $s_{i}\in\tau^{S}$, to compute the loss as follows:
+where $\mathbf{p}_{i}=[p_{1},p_{2},...,p_{L}]^{T}$ with $\sum\limits_{p_{k}\in\mathbf{p}_{i}}p_{k}=1$ representing the segment prediction likelihood.
+The predictions are used, together with the respective ground-truth segment label $s_{i}\in\tau^{S}$, to compute the loss as follows:
 $\mathcal{L}_{S}=\sum_{i}NLLLoss(\mathbf{p}_{i},s_{i}).$ (4)
@@
 IV EXPERIMENTAL EVALUATION
-To support the experiments under realistic and challenging cross-horticultural domains, we introduce a new dataset comprising four sequences from distinct horticultural crops, including detailed annotations for ground truth evaluation.
-This section outlines the evaluation protocol, implementation details, and training procedures, including preprocessing steps and the architecture of the proposed approach.
-Additionally, quantitative metrics and qualitative retrieval results are reported, showcasing the effectiveness of the approach.
-Furthermore, we conduct a runtime analysis to assess the computational efficiency of the PointNetPGAP.
-Figure 3: 2D and 3D visualizations of the sequences’ paths. The 2D representation shows the individual segments, while the 3D representation outlines the overlapping paths.
+This section outlines the evaluation protocol, implementation details, and training procedures of the proposed approach.
+Both quantitative and qualitative retrieval results are presented to support the proposed contributions.
+Additionally, a runtime analysis is performed to assess the computational efficiency of PointNetPGAP.
 IV-A The HORTO-3DLM Dataset
@@
 For example, an orchard sequence recorded in November 2022 is named ON22.
-The columns have the following meaning: name of the sequence (Sequence), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
-Sequence M Y Total Scans N.Rs N.Segs D [m] Plantation Type
+The columns have the following meaning: name of the sequence (Seq), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
+Seq. M Y NºScans N.Rs N.Segs D [m] Plantation Type
 Orchards
@@
 SJ23 June 2023 6389 3 5 742 tunnels
-TABLE II: Training and testing statistics used in the experiments.
-Cross-validation
-Sequence Test Anchors Training Anchors Seq. Training Anchors
-OJ22 1797 1605 100
-OJ23 4801 1193 512
-ON22 4993 1210 495
-SJ23 3509 1107 598
-IV-B Methodology for Place Recognition
+TABLE II: Cross-Validation training and evaluation data. Each column represents the training and test data of the evaluated sequence (row 1)
+OJ22 OJ23 ON22 SJ23
+Nº Training Anchors ($r_{th}\leq 2m$) 1605 4801 1210 1107
+Nº Test Anchors ($r_{th}\leq 10\text{m}$) 1797 1193 4993 3509
+TABLE III: KITTI Odometry dataset and cross-domain training and evaluation data
+  KITTI Ondemtry dataset 00 02 05 06 08 Nºof Frames 4051 4661 2761 1101 4071 Nº of Anchors ($r_{th}\leq 10\text{m}$) 809 344 454 273 353 Cross Domain Training / Evaluation data Sequences Nº Anchors Training OJ22, OJ23, SJ23 ON22 1705 Validation 00 809 Test 02, 05, 06, 08 1424
+IV-B HORTO-3DLM Dataset for Place Recognition
 From the proposed HORTO-3DLM dataset described in Section IV-A, each sequence was split into various segments (S1, S2, …, S6) as illustrated in Fig.
@@
 Following the protocol defined in Section III-C1, $r_{th}$ was set to 2 m for the training data, which means that a positive pair exists whenever an anchor-positive pair is within a range of 2 m.
-Additionally, from all anchor-positive pairs, only the anchors that were at least 0.5 m apart were selected for training.
-For evaluation purposes, $r_{th}$ was set to 10 m, and all anchors were used.
-IV-C Evaluation Protocol
-The evaluation of the proposed approach is based on a k-fold cross-validation protocol, with $k=4$, which consists of training on three sequences and testing on the sequence that was left out.
-This is repeated four times to obtain results on all sequences.
-The dataset split for the cross-validation protocol is outlined in Table II.
-The entries have the following meaning: the first column (Sequence) identifies the sequence on which a given model was tested, meaning the model was trained on the remaining sequences; the second column (Test Anchors) refers to the number of anchors/loops with which the model was tested; the third column (Training Anchors) refers to the number of anchors/loops the model was trained with, corresponding to the sum of the training anchors of the remaining sequences; the fourth column (Seq.
-Training Anchors) refers to the number of training anchors each sequence contains individually.
-As for the performance metrics, the results obtained from the proposed experiments are reported using Recall for the top-k retrieved candidates (i.e., Recall@k).
+Additionally, from all training anchor-positive pairs, only the anchors that are at least 0.5 m apart were selected for training.
+IV-C Evaluation Protocols
+The proposed approach is evaluated based on two strategies: a 4-fold cross-validation protocol; and a cross-domain approach.
+Performance metrics are reported using Recall for the top-k retrieved candidates (i.e., Recall@k).
+IV-C1 Cross-Validation
+This protocol evaluates generalization capabilities within the same domain.
+The HORTO-3DLM dataset is split into four sequences, with models trained on three sequences and evaluated on the remaining one.
+This process is repeated four times to cover all sequences.
+The dataset split is outlined in Table II.
+A retrieved candidate is considered a true positive if it is within an $10$ m radius of the query and belongs to the same segment; otherwise, it is considered a false positive.
+IV-C2 Cross-Domain
+This protocol assesses generalization capabilities across domains, specifically from horticultural to urban environments.
+The models are trained on the HORTO-3DLM dataset and evaluated on the KITTI Odometry dataset [8].
+The data split regarding this evaluation strategy is presented in Table III.
+A retrieved candidate is considered a true positive if it is within an $10$ m radius of the query (no segments are considered); otherwise, it is considered a false positive.
 IV-D Training and Implementation Details
@@
 The SOTA models were implemented based on the original code.
-All models were trained with 200 epochs on an NVIDIA GeForce RTX 3090 GPU, using the closest positive and 20 negatives ($m=20$) for each anchor.
+All models were trained with 200 epochs, using early stopping to obtain the best performance on the validation set.
+All models were trained on an NVIDIA GeForce RTX 3090 GPU, using the closest positive and 20 negatives ($m=20$) for each anchor.
 The margin value of the contrastive loss was set to 0.5 ($m=0.5$), and the model parameters were optimized using the AdamW optimizer with a learning rate ($L_{r}$) of 0.0001 and a weight decay ($W_{d}$) of 0.0005.
 Moreover, all proposed experiments were conducted on Python 3.8, and PyTorch with CUDA 11.6.
+TABLE IV: Recall@1 performance of the cross-validation protocol (same domain).
+Reporting the results of the models trained: (without SLC)/(with SLC).
+Bolded scores indicate the highest value between models trained without and with SLC, while underlined scores represent the highest score in each column (i.e., the best-performing model for each sequence).
+  Model OJ22 OJ23 ON22 SJ23 MEAN OverlapTransformer [5] 0.084/0.075-0.009 0.153/0.164+0.011 0.221/0.194-0.027 0.084/0.073-0.011 0.136/0.127-0.009 LOGG3D [7] 0.357/0.394+0.037 0.309/0.329+0.020 0.412/0.470+0.058 0.254/0.285+0.031 0.333/0.370+0.037 PointNetVLAD [6] 0.560/0.659+0.099 0.485/0.502+0.017 0.611/0.625+0.014 0.676/0.717+0.041 0.583/0.630+0.043 PNPGAP(ours) 0.873/0.880+0.007 0.643/0.688+0.045 0.685/0.734+0.049 0.678/0.712+0.034 0.719/0.754+0.034
 IV-E Retrieval Evaluation and Discussion
-This section presents and discusses the empirical results.
-Firstly, we conduct an ablation study on the proposed PointNetPGAP model, delineating the contribution of each aggregator to the performance.
-Secondly, we benchmark PointNetPGAP, both trained with and without the SLC model, against SOTA models: PointNetVLAD [9], LOGG3D-Net [10], and OverlapTransformer [8].
-Each of these models adopts distinct strategies for handling scans: PointNetVLAD processes scans directly, OverlapTransformer processes scans as range images, and LOGG3D-Net relies on a voxel-based representation and network.
-The main results are presented in Table III and Fig.
-4, while results at the segment-level are presented in Fig.
-3.
-Furthermore, a qualitative assessment is outlined in Fig.
-6.
-Finally, the efficiency of the models is evaluated through runtime analysis.
-Throughout this section, PointNetPGAP will be abbreviated as PNPGAP, and PNPGAP-SLC when trained without and with the SLC model, respectively.
+This section presents and discusses the empirical results of the work.
+First, an ablation study is conducted on the proposed PointNetPGAP model to showcase the contribution of each aggregator to the overall performance.
+Subsequently, PointNetPGAP is benchmarked within the same domain, both with and without the SLC model, against SOTA models: PointNetVLAD [6], LOGG3D-Net [7], and OverlapTransformer [5].
+These models employ different strategies for processing LiDAR scans: PointNetVLAD processes scans directly, OverlapTransformer processes scans as range images, and LOGG3D-Net utilizes a voxel-based representation.
+Finally, the generalization capabilities of the proposed approach across domains are evaluated.
+Throughout this section, PointNetPGAP is abbreviated as PNPGAP, and PNPGAP-SLC when trained with the SLC model.
+TABLE V: Ablation study of PNPGAP, where GAP represents the global average pooling aggregator and PFI represents the aggregator based on pairwise feature interactions.
+The bolded scores represent the highest values in each column.
+  GAP PFI Recall@1 Recall@1% OJ22 OJ23 ON22 SJ24 MEAN OJ22 OJ23 ON22 SJ24 MEAN 0.819 0.589 0.687 0.714 0.701 0.928 0.909 0.962 0.975 0.944 0.827 0.613 0.685 0.713 0.710 0.962 0.925 0.950 0.988 0.956 0.880 0.688 0.734 0.723 0.756 0.943 0.938 0.985 0.983 0.962
 IV-E1 Ablation Study
-The results of the ablation study are reported in Table IV.
-For this study, we report the Recall@1 and Recall@1% of each aggregation approach separately and combined.
-The results indicate that combining both aggregators yields higher performance compared to each aggregator individually, improving performance by 0.6 percentage points (pp) for retrieving the Top-1% candidates and 4.7 pp for retrieving the Top-1 candidate.
-These results suggest that the real impact of combining these aggregation approaches is on modeling the closest neighbor.
-TABLE III: Recall@1 performance of the models trained: (without SLC)/(with SLC).
-The bolded scores represent the highest value between trained without and with SLC, while the underling indicates the highest score in each column.
-  Model OJ22 OJ23 ON22 SJ23 MEAN OverlapTransformer [8] 0.084/0.075-0.009 0.153/0.164+0.011 0.221/0.194-0.027 0.084/0.073-0.011 0.136/0.127-0.009 LOGG3D [10] 0.357/0.394+0.037 0.309/0.329+0.020 0.412/0.470+0.058 0.254/0.285+0.031 0.333/0.370+0.037 PointNetVLAD [9] 0.560/0.659+0.099 0.485/0.502+0.017 0.611/0.625+0.014 0.676/0.717+0.041 0.583/0.630+0.043 PNPGAP(ours) 0.873/0.880+0.007 0.643/0.688+0.045 0.685/0.734+0.049 0.678/0.712+0.034 0.719/0.754+0.034
+The ablation study is conducted following the cross-validation protocol outlined in Section IV-C1, with results presented in Table V.
+For this analysis, we report Recall@1 and Recall@1% for each aggregation approach, both individually and combined.
+The results suggest that combining both aggregators results in superior performance compared to using each aggregator separately, with an improvement of 0.6 percentage points (pp) in retrieving the Top-1% candidates and 4.7 pp for retrieving the Top-1 candidate.
+These findings indicate that the proposed aggregator significantly enhances the model’s ability to identify the nearest neighbor.
 OJ22 Refer to caption
@@
 Figure 5: Recall@1 performance at the segment-level for the four sequences. For more details on the segments, please see Fig. 3.
-TABLE IV: Ablation study of PNPGAP, where GAP represents the global average pooling aggregator and PFI represents the aggregator based on pairwise feature interactions.
-The bolded scores represent the highest values in each column.
-  GAP PFI Recall@1 Recall@1% OJ22 OJ23 ON22 SJ24 MEAN OJ22 OJ23 ON22 SJ24 MEAN 0.819 0.589 0.687 0.714 0.701 0.928 0.909 0.962 0.975 0.944 0.827 0.613 0.685 0.713 0.710 0.962 0.925 0.950 0.988 0.956 0.880 0.688 0.734 0.723 0.756 0.943 0.938 0.985 0.983 0.962
-IV-E2 Comparison to SOTA
-PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD (the second-best model) by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the top-1 candidate.
-This is also evident in Fig.
-6, where a qualitative assessment is shown on the paths, displaying the true positive predictions of each model when retrieving the top-1 candidate.
+IV-E2 Cross-Validation (same domain)
+The results of this experiment are presented in Table IV.
+PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the Top-1 candidate.This superiority is further illustrated in Fig.
+6, where a qualitative assessment shows the true positive predictions of each model for the Top-1 candidate retrieval along the paths.
+Given that PNPGAP and PointNetVLAD share the same backbone, these results suggest that the proposed aggregator is more effective in horticultural environments, particularly in handling sparse scans, than other first- and second-order aggregators.
 When analyzing the performance of retrieving the Top-25 candidates (Fig.
@@
 Meanwhile, OverlapTransformer, which relies on range images and on a CNN-based network for feature extraction, shows poor performance across all sequences.
-The inadequate performance of OverlapTransformer is also apparent in Fig.
+The low performance of OverlapTransformer is also apparent in Fig.
 6, where in some segments the model fails to predict any loops correctly.
 This suggests the unsuitability of this model for agricultural environments.
-At the segment-level, most models perform best in segments 3 and 4, representing the extremities of the crops.
+When adding the SLC model at training time (as shown in Table IV), the results indicate that this additional training signal enhances performance across all sequences and models (except for OverlapTransformer), with an average performance increase of 3.7 pp for LOGG3D, 4.3 pp for PointNetVLAD, and 3.34 pp for the proposed PNPGAP model.
+At the segment-level (see Fig.5), most models (without SLC) achieve their highest performance in segments 3 and 4, which represent the extremities of the crops.
 As illustrated in Fig.
-1, these segments are distinctive, while segments representing the rows exhibit higher similarities and thus create ambiguities among neighboring rows, resulting in lower retrieval performance.
-Figure 6: True positive predictions for retrieving the top-1 candidate.
-IV-E3 Training with the SLC Model
-The results when adding the SLC model at training time (presented in Table III) indicate that this additional training signal improves the performance throughout all sequences and models (except OverlapTransformer), with an average performance increase of 3.7 pp for LOGG3D, 4.3 pp for PointNetVLAD, and 3.34 pp for the proposed PNPGAP model.
-These findings can be further analyzed in more detail in Fig.
-3, where a comparison between PNPGAP and PNPGAP-SLC is made at the segment-level.
-The results show that, while the performance in segments with high performance remains mostly unchanged, segments such as those corresponding to rows (i.e., 1, 2, and 5) increase performance, indicating that the SLC model produces better descriptors that are less susceptible to ambiguities.
+1,these segments are more distinctive.
+In contrast, segments representing the rows exhibit greater similarity, creating ambiguities among adjacent rows, which leads to lower retrieval performance.
+However, when the SLC model is incorporated, the results at the segment level indicate a notable performance increase, particularly in the segments corresponding to rows (i.e., 0, 1, 2, and 5).
+This improvement suggests that the SLC model produces more robust descriptors that are less susceptible to ambiguities.
+Figure 6: True positive predictions of the top-1 retrieved candidates.
+TABLE VI: Recall@1 of the cross-domain evaluation: (without SLC)/(with SLC).
+The bolded scores represent the highest value between trained without and with SLC, while the underlined scores indicate the highest score in each column.
+  Validation Testing 00 02 05 06 08 MEAN Overlap Transformer [5] 0.247/0.278+0.031 0.192/0.253+0.061 0.264/0.361+0.097 0.227/0.355+0.128 0.023/0.040+0.017 0.176/0.252+0.076 LOGG3D [7] 0.043/0.083+0.040 0.084/0.116+0.032 0.108/0.176+0.068 0.271/0.209-0.062 0.043/0.043+0.000 0.127/0.138+0.010 PointNetVLAD [6] 0.841/0.845+0.004 0.442/0.451+0.009 0.663/0.738+0.075 0.934/0.978+0.044 0.123/0.234+0.111 0.541/0.600+0.060 PNPGAP(ours) 0.826/0.841+0.015 0.395/0.448+0.053 0.683/0.659-0.024 0.938/0.971+0.033 0.248/0.308+0.060 0.566/0.597+0.031
+IV-E3 Cross Domain
+The results of the cross-domain experiment are presented in Table VI.
+In this experiment, the performance difference between PNPGAP and PointNetVLAD is smaller, with PNPGAP outperforming by 3.3 pp.
+This difference is primarily attributable to PNPGAP’s superior performance on sequence 08, the most challenging sequence, where most revisits occur from the reverse direction.
+This finding suggests that, while the proposed aggregation approach provides less of an advantage in urban environments compared to horticultural settings, it still demonstrates superior rotation-invariant properties relative to other SOTA models, particularly PointNetVLAD.
+When incorporating the SLC model during training, the performance improvements observed across models are consistent with those seen within the same domain, with the exception of OverlapTransformer.
+Notably, OverlapTransformer’s performance is significantly higher in the cross-domain experiment compared to its results on the HORTO-3DLM dataset, despite being trained on the same dataset in both cases.
+We hypothesize that this performance gain is primarily due to OverlapTransformer’s input representation, which appears to be better suited to the structured nature of urban environments.
 IV-F Runtime Analysis
-TABLE V: The average runtime per batch, where the batch size is 20 scans.
+Processing capacity and model size are critical factors for real-world implementation, particularly in resource-constrained robotics applications.
+The runtime experiment results, presented in Table VII, are obtained using the hardware detailed in Section IV-D.
+The reported times, measured in milliseconds (ms), represent the duration required by each model to process a batch of 20 scans.
+Retrieval duration was excluded from this experiment, as it remains consistent across all models.
+Model size is expressed by the number of parameters, in millions (M).
+PNPGAP is notably compact, with 0.4 million parameters, and processes a batch in an average of 4.4 ms.
+Comparative analysis with other SOTA models reveals that PNPGAP is approximately 5% the size of LOGG3D-Net and 31 times faster.
+Compared to PointNetVLAD, PNPGAP is only 2% of its size and 6 times faster.
+These results underscore the computational efficiency of the proposed PNPGAP model.
+TABLE VII: The average runtime per batch, where the batch size is 20 scans.
 Param. [M] OJ22 [ms] OJ23 [ms] ON22 [ms] SJ23 [ms] MEAN [ms]
@@
 PNPGAP (ours) 0.4 3.3 3.5 3.3 3.4 3.38
-The processing capacity and model size are crucial factors for real-world implementation, particularly in resource-constrained robotics applications.
-Throughout the runtime experiment, the results presented in Table V were obtained using identical hardware specifications as detailed in Section IV-D: an AMD Ryzen 9 5900X 12-Core CPU with 64 GB RAM and a GeForce RTX 3090 24GB GPU.
-The reported outcomes specifically denote the duration, expressed in milliseconds (ms), that each model required for processing a batch of 20 scans.
-The retrieval duration was omitted from this experiment because it remains uniform across all models.
-The size of the models is reported through the count of parameters, denominated in millions (M).
-PNPGAP emerges as one of the smaller models, comprising 0.4 million parameters and demanding on average 4.4 ms per batch.
-A comparative analysis against other SOTA models reveals that PNPGAP has approximately 5% of the size of LOGG3D-Net and is 31 times faster.
-Furthermore, in comparison to PointNetVLAD, PNPGAP is only 2% of its size and is 6 times faster.
-These outcomes highlight the computational efficiency of the proposed PNPGAP model.
 V CONCLUSIONS
-In this letter, we addressed the 3D LiDAR place recognition problem in horticultural settings by introducing a novel aggregation approach that fuses two statistic-inspired aggregators into a model called PointNetPGAP; an SLC model to handle inter-row descriptor ambiguities; and a novel dataset, HORTO-3DLM, which includes four sequences: three from orchards and one from a strawberry plantation.
-The experimental evaluation was conducted on the HORTO-3DLM dataset, and the proposed PointNetPGAP model was benchmarked against SOTA models.
-An ablation study demonstrated the superiority of the combined aggregators, while the results at both the sequence and segment levels show the proposed model outperforming SOTA models while being more efficient.
-Regarding the proposed training approach, all models were trained with and without the SLC model.
-The results showed that, except for one model, incorporating the SLC model during training led to an increase in performance.
-This improvement was primarily observed in segments with high inter-row similarity.
-This suggests that providing a context-specific training signal, as offered by the SLC model, may be crucial for effectively training place recognition models in environments where LiDAR scans overlap significantly, even when originating from different rows.
+In this letter, we addressed the challenge of 3D LiDAR-based place recognition in horticultural environments by introducing the PointNetPGAP model, which integrates two statistically-inspired aggregators, along with an SLC model designed to mitigate inter-row descriptor ambiguities.
+We also presented the HORTO-3DLM dataset, comprising sequences from orchards and a strawberry plantation.
+Experimental results obtained from the HORTO-3DLM dataset, using a cross-validation protocol, as well as from the KITTI Odometry dataset, suggest that PointNetPGAP outperforms SOTA models, as evidenced by an ablation study and benchmarks at both the sequence and segment levels.
+Furthermore, incorporating the SLC model during training consistently improved performance, particularly in segments with high inter-row ambiguity, highlighting the critical role of context-specific training signals.
 ACKNOWLEDGMENTS

### v4

@@
 The dataset and the code will be made publicly available at https://github.com/Cybonic/PointNetPGAP-SLC.git
+Index Terms:
+Deep Learning Methods; Localization; Agricultural Automation
 Figure 1: Geolocation, 3D maps and recording setup of the HORTO-3DLM dataset.
 Four data sequences were recorded over two years at different locations in the UK, using a Husky mobile platform equipped with a 32-beam Velodyne sensor and ZED-F9P RTK GPS/GNSS system.
-00footnotetext: 1T.Barros, L.Garrote, P.Conde, C.Premebida and U.J.Nunes are with the University of Coimbra, Institute of Systems and Robotics, Department of Electrical and Computer Engineering, Portugal.
+00footnotetext: This work has been supported by the project GreenBotics (ref.
+PTDC/EEI-ROB/2459/2021), funded by Fundação para a Ciência e a Tecnologia (FCT), Portugal, and by Innovate UK under the Grant No.
+10072930 and 10073332.
+It was also partially supported by FCT through grant UIDB/00048/2020 and under the PhD grant with reference 2021.06492.BD.00footnotetext: 1T.Barros, L.Garrote, P.Conde, C.Premebida and U.J.Nunes are with the University of Coimbra, Institute of Systems and Robotics, Department of Electrical and Computer Engineering, Portugal.
 E-mails: {tiagobarros, garrote, pedro.conde,  cpremebida, urbano}@isr.uc.pt00footnotetext: 2M.J.Coombes and C.Liu are with the Dept.
@@
 In recent years, significant advancements have been made in 3D LiDAR-based place recognition, driven primarily by the autonomous driving community with the aim of achieving long-term localization [1, 2].
-While initial efforts have been made to adapt these methods for agricultural environments [3, 4], several challenges unique to these settings remain largely unresolved.
+Although initial efforts have been made to adapt these methods for agricultural environments [3, 4], several challenges unique to these settings remain largely unresolved.
 Unlike urban environments, where LiDAR scans capture well-defined geometries (e.g., corners, planes), horticultural environments produce scans with sparse and poorly defined features.
@@
 Although its impact is reduced in urban settings, it nonetheless exhibits superior rotation invariance in urban environments compared to SOTA models.
-Moreover, when models are trained with the SLC approach, results at both the sequence and segment levels indicate improved , particularly by improving retrieval accuracy in segments with higher ambiguity.
+Moreover, when models are trained with the SLC approach, results at both the sequence and segment levels indicate improved retrieval performance, particularly in segments with higher ambiguity.
 Figure 2: Top: The retrieval framework uses PointNetPGAP to generate global descriptors through the following stages: T1) PointNet processes an input scan $P$ to extract local features $F$; T2) These features are aggregated into a global descriptor $D$ by combining global average representation with pairwise feature interactions; T3) The global descriptor is then used to query the database and retrieve the top-$k$ most similar places.
@@
 In the context of 3D LiDAR-based methods, although hand-crafted descriptors are still being used [9], Deep Learning (DL) techniques have emerged as the most prevalent choice for place modeling [1].
-DL-based place recognition models can be broadly split into two main components: a local feature extraction module that maps the input scan to a local feature space, and a global feature aggregation module that aggregates the local features into a global descriptor.
+DL-based place recognition models can be broadly split into two main components: a local feature extraction module that maps the input scan to a local feature space, and a feature aggregation module that aggregates the local features into a global descriptor.
 As for the feature extraction module, approaches can be categorized based on how the point clouds are processed: those that extract features directly from point clouds, such as PointNetVLAD [6], and those that resort to proxy representations, such as voxels [10, 11], polar coordinates [12], or depth range images [13, 5].
 Recent works such as MinkLoc3D [11], LCDNet [14], and LOGG3D-Net [7] tend to primarily use voxel-based approaches, employing sparse convolutions such as PV-RCNN [15] or 4D Spatio-Temporal ConvNets [16], which have shown higher feature extraction capabilities in urban-like environments.
-In this work, we resort to PointNet [17] due to its efficiency and lower computational demands (as demonstrated in Section IV-F).
+In this work, we resort to PointNet [17] due to its efficiency and lower computational demands (as demonstrated in Section IV-E).
 Regarding aggregation strategies, most 3D LiDAR place recognition frameworks rely on first-order aggregation approaches such as NetVLAD [6, 18, 14], which computes the residuals w.r.t.
@@
 Specifically, we combine the global average pooling aggregator with a pairwise feature interaction aggregator into a single descriptor.
-We compare our hypothesis with the SOTA models LOGG3D-Net [7], PointNetVLAD [6], and OverlapTransformer [5], and show through results on real-world datasets that the combination of these two statistical-based aggregators yields higher performance than each aggregator individually and also the SOTA models.
+We compare our hypothesis with the SOTA models LOGG3D-Net [7], PointNetVLAD [6], and OverlapTransformer [5], and show through empirical results on real-world datasets that the combination of these two statistical-based aggregators yields higher performance than each aggregator individually and also the SOTA models.
 II-B Datasets
@@
 This section introduces the PointNetPGAP and the proposed training regime, where a traditional contrastive learning approach is combined with the proposed SLC model.
-Figure 2 outlines both PointNetPGAP in a retrieval framework (Top) and PointNetPGAP in the proposed training scheme (Bottom), with the combination of both the contrastive and the Negative Log-Likelihood (NLL) losses.
+Figure 2 outlines both PointNetPGAP in a retrieval framework (Top) and PointNetPGAP in the proposed training scheme (Bottom).
 III-A Problem Formulation
@@
 III-B PointNetPGAP
-PointNetPGAP, here defined as a mapping function $\Theta$ (outlined in Fig.
-2 by T1, T2), can be decomposed into two functions: $\Theta\equiv\varphi\circ\phi$, where $\phi:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{n\times c}$ maps the input scan $P\in\mathbb{R}^{n\times 3}$ to hyper-dimensional features $F\in\mathbb{R}^{n\times c}$; and $\varphi:\mathbb{R}^{n\times c}\rightarrow\mathbb{R}^{d}$ aggregates the local features $F$ into a global descriptor $D\in\mathbb{R}^{d}$ with $d$ dimensions.
+PointNetPGAP, defined here as a mapping function $\Theta$ (outlined in Fig.
+2 by T1, T2), can be decomposed into two functions: $\Theta\equiv\varphi\circ\phi$, where $\phi:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{n\times c}$ maps the input scan $P\in\mathbb{R}^{n\times 3}$ to hyper-dimensional features $F\in\mathbb{R}^{n\times c}$ with $c$ dimensions; and $\varphi:\mathbb{R}^{n\times c}\rightarrow\mathbb{R}^{d}$ aggregates the local features $F$ into a global descriptor $D\in\mathbb{R}^{d}$ with $d$ dimensions.
 III-B1 Local Feature Extraction (T1)
@@
 III-C1 Training Data (B1)
-Place recognition methods traditionally resort to contrastive learning techniques, which learn to distinguish between similar (positive pairs) and dissimilar (negative pairs) data samples.
-In place recognition for urban environments, positive pairs are defined only by a fixed radius, which is a sufficient condition given the scale of the trajectories and revisits in these environments.
-However, to perform place recognition in crops with narrow rows, the radius is a necessary but not sufficient condition.
-The search conditions must guarantee that the search occurs along the row and does not transgress into neighboring rows.
-Dividing the crops into distinct segments and confining the search within each segment ensures the necessary conditions.
+Place recognition methods traditionally rely on contrastive learning techniques, which are designed to differentiate between similar (positive pairs) and dissimilar (negative pairs) data samples.
+In urban environments, positive pairs are typically defined by a fixed radius, which is a sufficient condition given the scale of the trajectories and revisits in such environments.
+However, in narrow-row crop environments, a fixed radius is a necessary but not sufficient condition for place recognition.
+Additional constraints are required to ensure that loops occur within the same row, rather than between neighboring rows.
+To address this, crops are divided into distinct segments, and loops must not only satisfy the radius condition but also originate from the same segment.
 Hence, in this work, a positive pair refers to a pair of scans that must meet the following constraints: (1) both scans must be within a $r_{th}$ range (in meters); (2) both scans must be from different revisits, meaning that the positive scan cannot be the immediate previous scan of the anchor; and (3) both scans must be from the same segment (see Fig.
@@
 $\mathcal{L}_{T}=\text{max}(d_{AP}-d_{AN}+m,0),$ (5)
-where $d_{AP}=\|D^{A}-D^{P}\|_{2}$ is the Euclidean distance between the anchor and the positive, $d_{AN}=\min\limits_{j}(\|D^{A}-\{D^{N}_{j}\}\|_{2})$ is the Euclidean distance between the anchor and the hardest negative i.e., the negative that is the closest w.r.t.
+where $d_{AP}=\|D^{A}-D^{P}\|_{2}$ is the Euclidean distance between the anchor and the positive, $d_{AN}=\min\limits_{j}(\|D^{A}-\{D^{N}_{j}\}\|_{2})$ is the Euclidean distance between the anchor and the hardest negative: i.e., the negative that is the closest w.r.t.
 $D^{A}$ in the descriptor space, while $m$ designates a margin value.
@@
 IV EXPERIMENTAL EVALUATION
-This section outlines the evaluation protocol, implementation details, and training procedures of the proposed approach.
+This section outlines the evaluation protocols, implementation details, and training procedures of the proposed approach.
 Both quantitative and qualitative retrieval results are presented to support the proposed contributions.
@@
 TABLE I: Summary of the four sequences.
-The various sequence names represent the first letter of plantation type, month and year of the recording.
-For example, an orchard sequence recorded in November 2022 is named ON22.
 The columns have the following meaning: name of the sequence (Seq), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
@@
   KITTI Ondemtry dataset 00 02 05 06 08 Nºof Frames 4051 4661 2761 1101 4071 Nº of Anchors ($r_{th}\leq 10\text{m}$) 809 344 454 273 353 Cross Domain Training / Evaluation data Sequences Nº Anchors Training OJ22, OJ23, SJ23 ON22 1705 Validation 00 809 Test 02, 05, 06, 08 1424
-IV-B HORTO-3DLM Dataset for Place Recognition
+IV-A1 HORTO-3DLM Dataset for Place Recognition
 From the proposed HORTO-3DLM dataset described in Section IV-A, each sequence was split into various segments (S1, S2, …, S6) as illustrated in Fig.
@@
 Additionally, from all training anchor-positive pairs, only the anchors that are at least 0.5 m apart were selected for training.
-IV-C Evaluation Protocols
+IV-B Evaluation Protocols
 The proposed approach is evaluated based on two strategies: a 4-fold cross-validation protocol; and a cross-domain approach.
 Performance metrics are reported using Recall for the top-k retrieved candidates (i.e., Recall@k).
-IV-C1 Cross-Validation
+IV-B1 Cross-Validation
 This protocol evaluates generalization capabilities within the same domain.
@@
 A retrieved candidate is considered a true positive if it is within an $10$ m radius of the query and belongs to the same segment; otherwise, it is considered a false positive.
-IV-C2 Cross-Domain
+IV-B2 Cross-Domain
 This protocol assesses generalization capabilities across domains, specifically from horticultural to urban environments.
@@
 A retrieved candidate is considered a true positive if it is within an $10$ m radius of the query (no segments are considered); otherwise, it is considered a false positive.
-IV-D Training and Implementation Details
+IV-C Training and Implementation Details
 All models were trained and evaluated under the same conditions.
@@
   Model OJ22 OJ23 ON22 SJ23 MEAN OverlapTransformer [5] 0.084/0.075-0.009 0.153/0.164+0.011 0.221/0.194-0.027 0.084/0.073-0.011 0.136/0.127-0.009 LOGG3D [7] 0.357/0.394+0.037 0.309/0.329+0.020 0.412/0.470+0.058 0.254/0.285+0.031 0.333/0.370+0.037 PointNetVLAD [6] 0.560/0.659+0.099 0.485/0.502+0.017 0.611/0.625+0.014 0.676/0.717+0.041 0.583/0.630+0.043 PNPGAP(ours) 0.873/0.880+0.007 0.643/0.688+0.045 0.685/0.734+0.049 0.678/0.712+0.034 0.719/0.754+0.034
-IV-E Retrieval Evaluation and Discussion
+IV-D Retrieval Evaluation and Discussion
 This section presents and discusses the empirical results of the work.
@@
   GAP PFI Recall@1 Recall@1% OJ22 OJ23 ON22 SJ24 MEAN OJ22 OJ23 ON22 SJ24 MEAN 0.819 0.589 0.687 0.714 0.701 0.928 0.909 0.962 0.975 0.944 0.827 0.613 0.685 0.713 0.710 0.962 0.925 0.950 0.988 0.956 0.880 0.688 0.734 0.723 0.756 0.943 0.938 0.985 0.983 0.962
-IV-E1 Ablation Study
-The ablation study is conducted following the cross-validation protocol outlined in Section IV-C1, with results presented in Table V.
+IV-D1 Ablation Study
+The ablation study is conducted following the cross-validation protocol outlined in Section IV-B1, with results presented in Table V.
 For this analysis, we report Recall@1 and Recall@1% for each aggregation approach, both individually and combined.
@@
 Figure 5: Recall@1 performance at the segment-level for the four sequences. For more details on the segments, please see Fig. 3.
-IV-E2 Cross-Validation (same domain)
+IV-D2 Cross-Validation (same domain)
 The results of this experiment are presented in Table IV.
-PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the Top-1 candidate.This superiority is further illustrated in Fig.
+PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the Top-1 candidate.This superiority is illustrated further in Fig.
 6, where a qualitative assessment shows the true positive predictions of each model for the Top-1 candidate retrieval along the paths.
@@
 When adding the SLC model at training time (as shown in Table IV), the results indicate that this additional training signal enhances performance across all sequences and models (except for OverlapTransformer), with an average performance increase of 3.7 pp for LOGG3D, 4.3 pp for PointNetVLAD, and 3.34 pp for the proposed PNPGAP model.
-At the segment-level (see Fig.5), most models (without SLC) achieve their highest performance in segments 3 and 4, which represent the extremities of the crops.
+At the segment-level (see Fig.
+5), most models (without SLC) achieve their highest performance in segments 3 and 4, which represent the extremities of the crops.
 As illustrated in Fig.
@@
 In contrast, segments representing the rows exhibit greater similarity, creating ambiguities among adjacent rows, which leads to lower retrieval performance.
-However, when the SLC model is incorporated, the results at the segment level indicate a notable performance increase, particularly in the segments corresponding to rows (i.e., 0, 1, 2, and 5).
+However, when the SLC model is incorporated, the results at the segment level indicate a notable performance increase, particularly in the segments corresponding to rows.
 This improvement suggests that the SLC model produces more robust descriptors that are less susceptible to ambiguities.
@@
   Validation Testing 00 02 05 06 08 MEAN Overlap Transformer [5] 0.247/0.278+0.031 0.192/0.253+0.061 0.264/0.361+0.097 0.227/0.355+0.128 0.023/0.040+0.017 0.176/0.252+0.076 LOGG3D [7] 0.043/0.083+0.040 0.084/0.116+0.032 0.108/0.176+0.068 0.271/0.209-0.062 0.043/0.043+0.000 0.127/0.138+0.010 PointNetVLAD [6] 0.841/0.845+0.004 0.442/0.451+0.009 0.663/0.738+0.075 0.934/0.978+0.044 0.123/0.234+0.111 0.541/0.600+0.060 PNPGAP(ours) 0.826/0.841+0.015 0.395/0.448+0.053 0.683/0.659-0.024 0.938/0.971+0.033 0.248/0.308+0.060 0.566/0.597+0.031
-IV-E3 Cross Domain
+IV-D3 Cross-Domain
 The results of the cross-domain experiment are presented in Table VI.
@@
 We hypothesize that this performance gain is primarily due to OverlapTransformer’s input representation, which appears to be better suited to the structured nature of urban environments.
-IV-F Runtime Analysis
+IV-E Runtime Analysis
 Processing capacity and model size are critical factors for real-world implementation, particularly in resource-constrained robotics applications.
-The runtime experiment results, presented in Table VII, are obtained using the hardware detailed in Section IV-D.
-The reported times, measured in milliseconds (ms), represent the duration required by each model to process a batch of 20 scans.
+The runtime experiment results, presented in Table VII, are obtained using the hardware detailed in Section IV-C.
+The reported times, measured in milliseconds (ms), represent the duration each model requires to process a batch of 20 scans.
 Retrieval duration was excluded from this experiment, as it remains consistent across all models.
@@
 ACKNOWLEDGMENTS
-This work has been supported by the project GreenBotics (ref.
-PTDC/EEI-ROB/2459/2021), funded by Fundação para a Ciência e a Tecnologia (FCT), Portugal, and by Innovate UK under the Grant No.
-10072930 and 10073332.
-It was also partially supported by FCT through grant UIDB/00048/2020 and under the PhD grant with reference 2021.06492.BD.
 The authors would also like to thank Dr Charles Whitfield at NIAB East Malling for facilitating orchard data collection campaigns.